Kitty19:

When Anna was born 16 years ago, her small town of Lewisville had a population of 15,000. The population has increased 11% each year since then. What is the current population of this town? a. The population of this town is approximately 68,925 b. The population of this town is approximately 79,663 c. The population of this town is approximately 17,324 d. The population of this town is approximately 240,000 Please select the best answer from the choices provided A B C D

rthrth:

The equation you need is 15000*1.11^16

Kitty19:

ok, thanks! :)

Mashupa:

\[x \left( t \right)=x _{0}\times \left(1+ r\right)t\]

Kitty19:

thank you! :)

Mashupa:

Initial Population X0 = 15000 Rate of Growth (%) (r) =11 of years (t) =16

Kitty19:

Thank you again!!! :)

Mashupa:

if you did the calculation right you should end up with B the correct answer

Kitty19:

Thank you again! XD XD XD XD

Mashupa:

15000 × (1 + 0.11)^16 =B =79663 sorry but i needed to clearify the equation for others.